# Generic motives and motivic cohomology of fields

### v1

@@ 3.3. Morphisms of generic motives and cycle modules
Let $(E,n)$ and $(F,m)$ be objects in ~$\tilde{\mathscr{E}}_{k}$ such that $F/k$ is finite. Then the following map:
$\bigoplus_{x\in\operatorname{Spec}(E\otimes_{k}F)}K_{m-n}^{M}(\kappa(x))$ $\rightarrow\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\})$
@@
where $\varphi_{x}:E\rightarrow\kappa(x)$ and $\psi_{x}:F\rightarrow\kappa(x)$ are the field extensions the morphism induced by the prime ideal $x\subset E\otimes_{k}F$.
@@
$\operatorname{Hom}_{\tilde{\mathscr{E}}_{k}}\big{(}(E,n),(F,m)\big{)}\rightarrow\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\}\big{)}$
@@
In particular, only (D1), norm maps (D2) and multiplication by symbols (D3) are sufficient to describe all morphisms of generic motives in the preceding case.
@@
The proposition follows from the following computation:
$\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\}\big{)}$ $\stackrel{{\scriptstyle(1)}}{{\simeq}}\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}\big{(}\operatorname{\hat{M}}(E)\otimes\operatorname{\hat{M}}(F),\mathbbm{1}\{m-n\}\big{)}$
@@
where the first two isomorphisms follow as in 3.1.8, using in particular that $\operatorname{\hat{M}}(F)=\operatorname{M}(F)$ is rigid, with strong dual $\operatorname{M}(F)^{*}=\operatorname{M}(F)$, and the last one follows from (2.2.1.d). The assertion about the description of the isomorphism follow from the construction of data (D1), (D2), (D3) recalled above. ∎
The presentation of the category ~$\tilde{\mathscr{E}}_{k}$ given in 6.0.4 implies a vanishing of morphisms in ~$\tilde{\mathscr{E}}_{k}$ in certain degree. It is notable that this vanishing does hold in the category of generic motives, at least after inverting the characteristic exponent $e$ of $k$.
@@
For any function field $E$ and $F$ over $k$, and any couple of integers $(n,m)\in\mathbb{Z}^{2}$ such that $n-m>\operatorname{degtr}_{k}(F)$, one has:
$\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\})_{e}=0.$
@@
One can assume that $n=0$ to simplify the notation. Put $d=\operatorname{degtr}_{k}(F)$. Let $X$ be a smooth model of $F$ over $k$. According to 3.1.2(2) and 3.1.11, one gets:
$\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(E),\operatorname{\hat{M}}(F)\{m\})=\underset{U\subset X}{\text{\rm``lim''}}\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)}(\operatorname{\hat{M}}(E),\operatorname{M}(U)\{m\})$
where the projective limit runs over the non-empty open subschemes of $X$. According to duality, which holds under our assumptions (see 2.1.3), one gets:
$\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)}$ $(\operatorname{\hat{M}}(E),\operatorname{M}(U)\{m\})_{e}$
$\simeq\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)}(\operatorname{\hat{M}}(E)\otimes\operatorname{M}^{c}(U),\mathbbm{1}(m+d)[m+2d])_{e}$
Moreover, the motive with compact support $\operatorname{M}^{c}(U)$ is effective. [^23] In characteristic 00, this is [VSF00, Chap. 5, Cor. 4.1.6]. In general, one uses resolution of singularities by alteration and the homotopy exact sequence: $\operatorname{M}^{c}(X-Z)\rightarrow\operatorname{M}^{c}(X)\rightarrow\operatorname{M}^{c}(Z)$ which holds in $\operatorname{DM}(k)$ thanks to the six functors formalism and [CD15]. Therefore, the cancellation theorem implies that the left hand-side vanishes whenever $m+d<0$ as required. ∎
4. Computations of generic motives
@@ 4.1. Rational curves
4.1.1.
Generic motives, as pro-$t$-projective objects relative to the homotopy $t$-structure (3.2.10), are huge. This can be seen from the following computation, that we recall from [Dég08b].
@@
induced by the following split short (homotopy) exact sequence in $\operatorname{DM^{(0)}}(k)$:
@@
the splitting $\phi=(\phi_{x})_{x\in\mathbb{A}^{1}_{F,\{0\}}}$ of the map $d$ being defined for a given point $x$ as follows:
@@
4.1.3.
As a corollary of the preceding proposition, one obtains a canonical splitting $s_{t}:\operatorname{\hat{M}}(F)\rightarrow\operatorname{\hat{M}}(F(t))$ of the map $\varphi^{*}:\operatorname{\hat{M}}(F(t))\rightarrow\operatorname{\hat{M}}(F)$, which is uniquely determined by the following commutative diagram:
@@
In fact, this splitting holds more generally.
@@
Consider a ring $R$ such that , $R=\mathbb{Q}$ if $\operatorname{char}(k)=0$, and $R=\mathbb{Z}[1/e]$ if $\operatorname{char}(k)=e>0$.
Then for any extension of function fields $\varphi:E\rightarrow L$, the map $\varphi^{*}:\operatorname{\hat{M}}(L)\rightarrow\operatorname{\hat{M}}(E)$ is a split epimorphism after tensoring with $R$.
@@
The case where $L/E$ is finite follows from the degree formula [^24] This formula is for example a consequence of (R2c) and the fact that the norm map associated to $\varphi$ is multiplication by $d$ on the degree 00 part of Milnor K-theory:
@@
where $d$ is the degree of $L/E$, invertible in the coefficient ring $R$ by our assumptions. In the general case, $L$ is a finite extension of a purely transcendental extension $E(t_{1},...,t_{n})$. We are reduced to the purely transcendental case which holds according to the paragraph preceding the statement. ∎
The following proposition contradicts the hope of 3.3.9: the functor $\rho$ of (3.3.1.a) cannot be full in general.
@@
$\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(k(t))\{n\}\big{)}\xleftarrow{\ \sim\ }\operatorname{K}^{M}_{n}(F)\oplus\prod_{\begin{subarray}{c}x\in\mathbb{A}^{1}_{k,(0)}\\ y\in\operatorname{Spec}(F\otimes_{k}\kappa_{x})\end{subarray}}\operatorname{K}^{M}_{n+1}(\kappa_{y})$
which to a symbol $\sigma\in\operatorname{K}^{M}_{n}(F)$ associates the morphism
@@
$\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(k(t)\{n\})$
$\simeq\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(F),\mathbbm{1}\{n\})\oplus\prod_{x\in\mathbb{A}^{1}_{k,(0)}}\underset{=\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(F)\otimes\operatorname{\hat{M}}(\kappa_{x}),\mathbbm{1}\{n\})}{\underbrace{\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}(\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(\kappa_{x})\{n\})}}$
One finally concludes using the isomorphism (3.3.5.a) and of course, as $\kappa_{x}/k$ is finite separable:
@@
The preceding proposition shows that the functor $\rho$ is not full, negatively answering a question in [Dég02, 9.3.2]. This comes from the fact generic motives are pro-objects. In particular, a correct formulation of the conjecture describing morphisms of generic motives could be formulated as follows.
Conjecture.
There exists a natural topology on morphisms of generic motives such that morphisms of type (Dn*), for $1\leq n\leq 4$, are topological generators. Relations among these topological generators are described by relations (Rm*) as spelled out in 6.0.1.
4.2. Positive genus curves
4.2.1.
@@ 4.2. Positive genus curves
Let $A$ be the albanese scheme of $\bar{C}$ — isomorphic to the dual abelian variety of the Jacobian variety $J(\bar{C},x_{0})$. There is a canonical map of $k$-schemes.
$C\rightarrow\bar{C}\rightarrow A.$
We let $\underline{A}$ be the sheaf of abelian groups represented by $A$ on $\mathrm{Sm}_{k}$. It is $\mathbb{A}^{1}$-invariant and admits transfers (see e.g. [Org04, SS03]), and therefore an object of $\operatorname{HI}(k)$. The preceding map induces a morphism of sheaves with transfers
@@
which therefore induces a map $\pi_{1}:\operatorname{M}(C)\rightarrow\underline{A}$ in $\operatorname{DM^{eff}}(k)$. We let $\pi_{0}=p_{*}:\operatorname{M}(C)\rightarrow\mathbbm{1}$ be the map induced by the projection.
According to a theorem of Suslin and Voevodsky (see [VSF00, Chap. 5, 3.4.2]), the following map of $\operatorname{DM^{eff}}(k)$
@@
is an isomorphism. One deduces the following computation of generic motives.
@@
Consider the above notation, and let $\pi_{i}^{K}$ be the composition of $\pi_{i}$ with the canonical map $\operatorname{\hat{M}}(K)\rightarrow\operatorname{M}(C)$.
Then we have a (homotopy) exact sequence in $\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)$:
@@
where we have denoted by $\partial_{x}$ the residue map (3.3.5(D4*)) associated with the valuation on $K$ corresponding to the closed point $x\in C$.
@@
it corresponds to the element $(x-x_{0})_{x\in C_{(0)}}$ where $(x-x_{0})$ is seen as a 00-cycle of degree 00 on $\bar{C}_{\kappa_{x}}$.
Therefore, as soon as the field $K$ is not rational over $k$, the boundary $\partial$ is non trival, and the above homotopy exact sequence is non-split. Therfore, the motive $\operatorname{\hat{M}}(K)$ is in some sense “maximally mixed”.
@@
Let $Z\subset C$ be a closed subscheme. As $k$ is perfect, $Z$ is smooth and one deduces a homotopy exact sequence in $\operatorname{DM^{eff}_{gm}}(k)$:
@@
Note that by additivity $\operatorname{M}(Z)=\prod_{x\in Z}\operatorname{\hat{M}}(\kappa_{x})$. Thus, taking the projective limit of the homotopy exact sequences over the open complement $U=C-Z$ yields a homotopy exact sequence in $\operatorname{pro-\mskip-3.0mu}\operatorname{DM^{eff}_{gm}}(k)$ which has the required form up to using the mentioned additivity and the isomorphism (4.2.1.a). The assertion about the boundary map $\partial$, which is nothing else than the projective limit of the Gysin morphisms $i_{Z}^{!}$, thus follows. ∎
4.2.3.
In [Dég11], we associate to any homotopy sheaf $F$ in $\operatorname{HI}(k)$ a homotopy module $F_{*}=\sigma^{\infty}(F)$ (see also 2.1.2) with the formula for $n\geq 0$:
@@
where the tensor product and internal Hom are taken with respect to the canonical monoidal structure on $\operatorname{HI}(k)$. [^25] Recall that $F_{-n}$ is also obtained by iterating Voevodsky’s $(-1)$-construction; see loc. cit. 1.13.
This construction can be described more precisely when $F=\underline{A}$ is the homotopy sheaf associated with an abelian variety, as recalled above. Let us recall this description due to Kahn and Yamazaki.
Proposition 4.2.4.
Let $A$ be an abelian variety over $k$, and $\underline{A}$ be the associated homotopy sheaf as defined above.
Then for any $n>0$ and any function field $F$, one has:
$\underline{A}_{-n}(F)$ $=0$
$\underline{A}_{n}(F)$ $=\operatorname{K}^{SM}_{n}(F;A)$
where $\operatorname{K}^{SM}_{n}(F;A)=K(F,A,\mathbb{G}_{m},\ldots,\mathbb{G}_{m})$ is Somekawa K-theory associated with the semi-abelian varieties given by one copy of $A$ and $n$ copies of $\mathbb{G}_{m}$ (see [Som90], or rather [KY13, Def. 5.1] for the correct sign).
The first computation implies $\underline{A}_{-1}=0$, which is equivalent to say that the sheaf $\underline{A}$ is a birational invariant: we refer the interested reader to [KS17] for more on birational (homotopy) sheaves.
@@
The first computation simply follows as $A(\mathbb{G}_{m,F})=A(F)$. The second computation follows from the main theorem of [KY13], Th. 1.12. ∎
We can now state the following corollary of 4.2.2.
@@
Consider the notation of the aforementioned proposition. Then for any integer $n\in\mathbb{Z}$, there exits a short exact sequence of abelian groups:
$\prod_{\begin{subarray}{c}x\in C_{(0)}\\ y\in\operatorname{Spec}(F\otimes_{k}\kappa_{x})\end{subarray}}\operatorname{K}^{M}_{n+1}(\kappa_{y})\rightarrow\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(K)\{n\}\big{)}\rightarrow\operatorname{K}^{M}_{n}(F)\oplus\operatorname{K}^{SM}_{n}(F;A)\rightarrow 0$
@@
$\operatorname{\hat{M}}(F)\xrightarrow{\psi_{y!}}\operatorname{\hat{M}}(\kappa_{y})\xrightarrow{\gamma_{\tau}}\operatorname{\hat{M}}(\kappa_{y})\{n+1\}\xrightarrow{\varphi_{y}^{*}}\operatorname{\hat{M}}(\kappa_{x})\{n+1\}\xrightarrow{\partial_{x}}\operatorname{\hat{M}}(K)\{n\}$
given the canonical morphisms $F\xrightarrow{\psi_{y}}\kappa_{y}\xleftarrow{\varphi_{y}}\kappa_{x}$, and the residue map $\partial_{x}$ associated with the valuation on $K$ corresponding to the point $x$ as defined in 3.3.5(D4*).
@@
$\operatorname{Hom}_{\operatorname{DM^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(F)\{1\},\operatorname{\hat{M}}(K)\big{)}\simeq\prod_{x\in C_{F,(0)}}\mathbb{Z}.$
@@
One applies the functor $\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)}(\operatorname{\hat{M}}(F),-)$ to the homotopy exact sequence of 4.2.2, and proceed as in the proof of 4.1.5 for the first term of the above exact sequence, use the preceding proposition to compute the third one, and then use the vanishing of motivic cohomology $\operatorname{H}_{\mathrm{M}}^{n+1,n}(L)$ for any field $L$ for the surjectivity assertion. ∎
We end-up this section by a remark about duality.
@@
Consider the notation of 4.2.1. Then the motive $\underline{A}$ is rigid in $\operatorname{DM}(k)$ and there exists a canonical isomorphism:
@@
According to 2.1.3, the motive $\operatorname{M}(C)$ is rigid, with dual $\operatorname{M}^{c}(C)(-1)[-2]$. The homotopy exact sequence of the motive with compact support

### v3

@@ 4.1. Initial remarks about generic motives
Let $(E,n)$ and $(F,m)$ be objects in ~$\tilde{\mathcal{E}}_{k}$ such that $F/k$ is finite. Then the following map is an isomorphism:
$\bigoplus_{x\in\operatorname{Spec}(E\otimes_{k}F)}K_{m-n}^{M}(\kappa(x))$ $\rightarrow\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\})$
@@
where $\varphi_{x}:E\rightarrow\kappa(x)$ and $\psi_{x}:F\rightarrow\kappa(x)$ are the canonical morphisms induced by the prime ideal $x\subset E\otimes_{k}F$.
@@
$\operatorname{Hom}_{\tilde{\mathcal{E}}_{k}}\big{(}(E,n),(F,m)\big{)}\rightarrow\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\}\big{)}$
@@
Thus, in the preceding case, only (D1), norm maps (D2) and multiplication by symbols (D3) are sufficient to describe all morphisms of generic motives.
@@
The proposition is obtained from the following computation:
$\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\}\big{)}$ $\stackrel{{\scriptstyle(1)}}{{\simeq}}\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}\big{(}\operatorname{\hat{M}}(E)\otimes\operatorname{\hat{M}}(F),\mathbbm{1}\{m-n\}\big{)}$
@@
where the first two isomorphisms follow as in 3.1.8, particularly using that $\operatorname{\hat{M}}(F)=\operatorname{M}(F)$ is rigid, with strong dual $\operatorname{M}(F)^{*}=\operatorname{M}(F)$, and the last one follows from (2.2.1.d). The assertion regarding the description of the isomorphism follows from the construction of data (D1), (D2), (D3) as recalled above. ∎
The presentation of the category ~$\tilde{\mathcal{E}}_{k}$ given in 7.0.4 implies a vanishing of morphisms in ~$\tilde{\mathcal{E}}_{k}$ in certain degrees. It is notable that this vanishing also holds in the category of generic motives, at least after inverting the characteristic exponent $e$ of $k$.
@@
For all function fields $E$ and $F$ over $k$, and for all pairs of integers $(n,m)\in\mathbb{Z}^{2}$ such that $n-m>\operatorname{trdeg}_{k}(F)$, one has:
$\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(E)\{n\},\operatorname{\hat{M}}(F)\{m\})[1/e]=0.$
@@
We may assume that $n=0$ to simplify the notation. Put $d=\operatorname{trdeg}_{k}(F)$. Let $X$ be a smooth model of $F$ over $k$. According to 3.1.2(2) and 3.2.1, one gets:
$\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(E),\operatorname{\hat{M}}(F)\{m\})=\underset{U\subset X}{\operatorname{lim}}\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)}(\operatorname{\hat{M}}(E),\operatorname{M}(U)\{m\})$
where the projective limit runs over the non-empty open subschemes of $X$. By applying duality, which holds under our assumptions (see 2.1.5), one gets:
$\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)}$ $(\operatorname{\hat{M}}(E),\operatorname{M}(U)\{m\})[1/e]$
$\simeq\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)}(\operatorname{\hat{M}}(E)\otimes\operatorname{M}^{c}(U),\mathbbm{1}(m+d)[m+2d])[1/e]$
Moreover, the motive with compact support $\operatorname{M}^{c}(U)$ is effective, as noted in 2.1.5. Thus, by the cancellation theorem, the left hand-side vanishes whenever $m+d<0$ as required. ∎
@@ 4.2. Rational curves
4.2.1.
Being (formal) projective limits, generic motives are inherently large objects. This is confirmed by the following computation, which we recall from [Dég08b].
@@
arising from the following split homotopy exact sequence in $\operatorname{\mathscr{D}M^{(0)}}(k)$:
@@
where the splitting $\phi=(\phi_{x})_{x\in\mathbb{A}^{1}_{F,\{0\}}}$ of the map $d$ is defined for each point $x$ as follows:
@@
4.2.3.
As a corollary of the preceding proposition, we obtain a canonical splitting $s_{t}:\operatorname{\hat{M}}(F)\rightarrow\operatorname{\hat{M}}(F(t))$ of the map $\varphi^{*}:\operatorname{\hat{M}}(F(t))\rightarrow\operatorname{\hat{M}}(F)$, which is uniquely determined by the following commutative diagram:
@@
Indeed, this splitting holds more generally.
@@
For any extension of function fields $\varphi:E\rightarrow L$, the map $\varphi^{*}:\operatorname{\hat{M}}(L)\rightarrow\operatorname{\hat{M}}(E)$ is a split epimorphism after $\mathbb{Q}$-localization.
@@
The case where $L/E$ is finite follows from the degree formula [^31] This formula is for example a consequence of (R2c) and the fact that the norm map associated with $\varphi$ acts as multiplication by $d$ on the degree 00 part of Milnor K-theory:
@@
where $d$ denotes the degree of $L/E$.
In the general case, $L$ is a finite extension of a purely transcendental extension $E(t_{1},\ldots,t_{n})$. We are reduced to the purely transcendental case which holds according to the paragraph preceding the statement. ∎
The following proposition contradicts the hope of 4.1.1: the functor $\rho$ of (3.3.1.a) cannot be full in general.
@@
$\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(k(t))\{n\}\big{)}\xleftarrow{\ \sim\ }\operatorname{K}^{M}_{n}(F)\oplus\prod_{\begin{subarray}{c}x\in\mathbb{A}^{1}_{k,(0)}\\ y\in\operatorname{Spec}(F\otimes_{k}\kappa_{x})\end{subarray}}\operatorname{K}^{M}_{n+1}(\kappa_{y})$
which to a symbol $\sigma\in\operatorname{K}^{M}_{n}(F)$ associates the morphism:
@@
$\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(k(t)\{n\})$
$\simeq\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(F),\mathbbm{1}\{n\})\oplus\prod_{x\in\mathbb{A}^{1}_{k,(0)}}\underset{=\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(F)\otimes\operatorname{\hat{M}}(\kappa_{x}),\mathbbm{1}\{n\})}{\underbrace{\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}(\operatorname{\hat{M}}(F),\operatorname{\hat{M}}(\kappa_{x})\{n\})}}$
One finally concludes using the isomorphism (3.3.4.a) and noting that $\kappa_{x}/k$ is finite separable:
@@
The preceding proposition shows that the functor $\rho$ is not full, negatively answering a question in [Dég02, 9.3.2]. This follows from the fact that generic motives are pro-objects. In particular, a more accurate formulation of the conjecture describing morphisms of generic motives could be formulated as follows.
Conjecture 4.2.7.
There exists a natural topology on morphisms of generic motives such that morphisms of type (Dn*), for $1\leq n\leq 4$, are topological generators. Relations among these topological generators are described by relations (Rm*) as spelled out in 7.0.1.
4.3. Curves of positive genus
4.3.1.
@@ 4.3. Curves of positive genus
The Jacobian $J$ of $\bar{C}/k$ is defined as the $k$-scheme representing the connected component of the Picard functor (see [GW23, Th. Def. 27.137]). On the other hand, the proper pointed $k$-scheme $(\bar{C},x_{0})$ admits a unique Albanese scheme $(A,\alpha:\bar{C}\rightarrow A)$ such that $\alpha$ maps $x_{0}$ to the zero element of $A$. In fact, letting $J^{\vee}$ be the dual abelian variety of $J$ over $k$, the base point $x_{0}$ defines a canonical map $\bar{C}\rightarrow J^{\vee}$, which satisfies the universal property of the Albanese scheme of $(\bar{C},x_{0})$, and we can just put $A=J^{\vee}$ (see [GW23, Rem. 27.225] for a clear account).
One considers the following composite map:
[…]
We let $\underline{A}$ be the sheaf of abelian groups represented by $A$ on $\mathrm{Sm}_{k}$. It is $\mathbb{A}^{1}$-invariant and admits transfers (see e.g. [Org04, SS03]), thus making it an object of $\operatorname{HI}(k)$. The preceding map induces a morphism of sheaves with transfers
@@
which gives us a map $\pi_{1}:\operatorname{M}(C)\rightarrow\underline{A}$ in $\operatorname{\mathscr{D}M^{eff}}(k)$. We define $\pi_{0}=p_{*}:\operatorname{M}(C)\rightarrow\mathbbm{1}$ as the map induced by the projection.
According to a theorem of Suslin and Voevodsky (see [VSF00, Chap. 5, 3.4.2]), the following map in $\operatorname{\mathscr{D}M^{eff}}(k)$
@@
is an isomorphism. From this, one deduces the following computation of generic motives.
@@
Consider the notation introduced above, and let $\pi_{i}^{K}$ be the composition of $\pi_{i}$ with the canonical map $\operatorname{\hat{M}}(K)\rightarrow\operatorname{M}(C)$.
Then we have a homotopy exact sequence (see 6.4.3) in $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)$:
@@
where we have denoted by $\partial_{x}$ the residue map (3.3.4(D4*)) associated with the valuation on $K$ corresponding to the closed point $x\in C$.
@@
it corresponds to the element $(x-x_{0})_{x\in C_{(0)}}$ where each $(x-x_{0})$ is seen as a 00-cycle of degree 00 on $\bar{C}_{\kappa_{x}}$.
Therefore, as soon as the field $K$ is not rational over $k$, the boundary map $\partial$ is non trivial, and the above homotopy exact sequence is non-split. This means in some sense that the motive $\operatorname{\hat{M}}(K)$ is “maximally mixed”.
@@
Let $Z\subset C$ be a closed subscheme. As $k$ is perfect, $Z$ is smooth and one deduces a homotopy exact sequence in $\operatorname{\mathscr{D}M^{eff}_{gm}}(k)$:
@@
Note that by additivity $\operatorname{M}(Z)=\prod_{x\in Z}\operatorname{\hat{M}}(\kappa_{x})$. Thus, taking the projective limit of these homotopy exact sequences over the open complement $U=C-Z$ yields a homotopy exact sequence in $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M^{eff}_{gm}}(k)$ which has the required form up to using the mentioned additivity and the isomorphism (4.3.1.a). The statement regarding the boundary map $\partial$, which is indeed the projective limit of the Gysin morphisms $i_{Z}^{!}$, thus follows. ∎
4.3.3.
In [Dég11], we associate to any homotopy sheaf $F$ over $k$ a homotopy module $F_{*}=\sigma^{\infty}(F)$ (see also 2.1.4) with the formula for $n\geq 0$:
@@
where the tensor product and internal Hom are taken with respect to the canonical monoidal structure on $\operatorname{HI}(k)$. [^32] Recall that $F_{-n}$ is also obtained by iterating Voevodsky’s $(-1)$-construction, recalled in (2.1.4.a). See [Dég11, 1.13] for details.
This construction can be described more precisely when $F=\underline{A}$ is the homotopy sheaf associated with an abelian variety (see 4.3.1). Let us recall this description due to Kahn and Yamazaki.
Theorem 4.3.4.
Let $A$ be an abelian variety over $k$, and $\underline{A}_{*}$ be the associated homotopy module as defined above.
Then for any $n>0$ and any function field $E$, one has:
$\underline{A}_{-n}(E)$ $=0$
$\underline{A}_{n}(E)$ $=\operatorname{K}^{SM}_{n}(E;A)$
where $\operatorname{K}^{SM}_{n}(E;A)=K(E,A,\mathbb{G}_{m},\ldots,\mathbb{G}_{m})$ is the Somekawa K-theory associated with the semi-abelian variety $A\times\mathbb{G}_{m}^{n}$ (see [Som90], or rather [KY13, Def. 5.1] for the correct sign).
@@
For the first isomorphism, we only need to prove that $\underline{A}_{-1}=0$. According to 3.1.10, it is only necessary to check that for any function field $E/k$, one has $\underline{A}_{-1}(E)=0$. This follows from the exact sequence (2.1.4.a) and the classical fact that the projection map $p$ induces an isomorphism $A(E)\xrightarrow{p^{*}}A(\mathbb{G}_{m,E})$, since $A$ is an abelian $k$-scheme.
The second computation follows from the main theorem of [KY13], stated as (1.1) in the introduction. ∎
Remark 4.3.5.
The vanishing $\underline{A}_{-1}=0$ is equivalent to say that the sheaf $\underline{A}$ is a “birational invariant”: we refer the interested reader to [KS17] for more on this notion (see [KS17, 4.1] for the case of abelian schemes).
We can now state the following corollary of 4.3.2.
@@
Consider the notations of the aforementioned proposition, and let $E$ be an arbitrary function field. Then for any integer $n\in\mathbb{Z}$, there exists a short exact sequence of abelian groups:
$\prod_{\begin{subarray}{c}x\in C_{(0)}\\ y\in\operatorname{Spec}(E\otimes_{k}\kappa_{x})\end{subarray}}\operatorname{K}^{M}_{n+1}(\kappa_{y})\rightarrow\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(E),\operatorname{\hat{M}}(K)\{n\}\big{)}\rightarrow\operatorname{K}^{M}_{n}(E)\oplus\operatorname{K}^{SM}_{n}(E;A)\rightarrow 0$
@@
$\operatorname{\hat{M}}(E)\xrightarrow{\psi_{y!}}\operatorname{\hat{M}}(\kappa_{y})\xrightarrow{\gamma_{\tau}}\operatorname{\hat{M}}(\kappa_{y})\{n+1\}\xrightarrow{\varphi_{y}^{*}}\operatorname{\hat{M}}(\kappa_{x})\{n+1\}\xrightarrow{\partial_{x}}\operatorname{\hat{M}}(K)\{n\}$
given the canonical morphisms $E\xrightarrow{\psi_{y}}\kappa_{y}\xleftarrow{\varphi_{y}}\kappa_{x}$, and the residue map $\partial_{x}$ associated with the valuation on $K$ corresponding to the point $x$ as defined in 3.3.4(D4*).
@@
$\operatorname{Hom}_{\operatorname{\mathscr{D}M^{(0)}}(k)}\big{(}\operatorname{\hat{M}}(E)\{1\},\operatorname{\hat{M}}(K)\big{)}\simeq\prod_{x\in C_{F,(0)}}\mathbb{Z}.$
@@
One applies the functor $\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)}(\operatorname{\hat{M}}(E),-)$ to the homotopy exact sequence of 4.3.2, and proceeds as in the proof of 4.2.5 for the first term of the above exact sequence, we use the preceding proposition to compute the third one, and then apply the vanishing of motivic cohomology $\operatorname{H}_{\mathrm{M}}^{n+1,n}(L)$ for any field $L$ for the surjectivity assertion. ∎
We end this section by a remark about duality.
@@
Consider the notation of 4.3.1. Then the motive $\underline{A}$ is rigid in $\operatorname{\mathscr{D}M}(k)$ and there exists a canonical isomorphism:
@@
According to 2.1.5, the motive $\operatorname{M}(C)$ is rigid, with dual given by $\operatorname{M}^{c}(C)(-1)[-2]$. The homotopy exact sequence of the motive with compact support